OpenStudy (anonymous):

The speed of sound V near the Earth's surface can be represented by this equation,V = 20 times the square root of t + 273 , where t is the temperature in degrees Celsius. Find the temperature when V = 380 meters per second. Round to the nearest whole number. A. 88 B. 90 C. 92 D. 94

OpenStudy (anonymous):

\[V=20\sqrt{t+273} is the problem for the question. \]

OpenStudy (anonymous):

is the problem for the question**

OpenStudy (anonymous):

\[380=20\sqrt{t+273}\]

OpenStudy (anonymous):

I don't get it.. :/

OpenStudy (anonymous):

\[380=20\sqrt{t+273}\] divide 20 on both sides \[19=\sqrt{t+273}\] square them so the radical cancel out \[19^2=(\sqrt{t+273})^2\] \[361=t+273\] 88=t